# Towards Global AI Inclusivity: A Large-Scale Multilingual Terminology Dataset

### v1

@@ Appendix C Additional Experiment Results
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 17.29 19.21 18.31
Bleu score std 22.88 24.37 24.26
Bleu 1gram precision mean 17.37 19.30 18.40
Bleu 1gram precision std 22.94 24.43 24.32
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 24.74 26.52 24.75
Bleu ref len std 12.69 13.49 12.72
Bleu pred len mean 24.59 26.35 24.60
Bleu pred len std 12.46 13.22 12.49
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 17.45 19.36 18.47
Chrf score std 23.08 24.54 24.43
Ter score mean 83.26 81.33 82.23
Ter score std 23.01 24.53 24.40
Chrfpp score mean 17.45 19.36 18.47
Chrfpp score std 23.08 24.54 24.43
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 26.26 27.23 27.06
Bleu score std 24.84 26.18 25.59
Bleu 1gram precision mean 26.51 27.50 27.32
Bleu 1gram precision std 24.89 26.23 25.65
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 28.70 29.35 28.76
Bleu ref len std 14.62 14.70 14.65
Bleu pred len mean 28.24 28.86 28.30
Bleu pred len std 13.91 13.97 13.96
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.03 0.03 0.03
Bleu ratio mean 0.99 0.99 0.99
Bleu ratio std 0.04 0.04 0.04
Chrf score mean 26.43 27.41 27.23
Chrf score std 24.92 26.25 25.66
Ter score mean 74.32 73.34 73.51
Ter score std 25.21 26.56 25.97
Chrfpp score mean 26.43 27.41 27.23
Chrfpp score std 24.92 26.25 25.66
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 32.64 34.24 33.30
Bleu score std 29.93 31.10 30.52
Bleu 1gram precision mean 32.69 34.30 33.34
Bleu 1gram precision std 29.94 31.12 30.53
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 29.97 30.02 29.94
Bleu ref len std 14.74 14.79 14.77
Bleu pred len mean 30.02 30.06 30.00
Bleu pred len std 14.73 14.74 14.77
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 32.76 34.36 33.42
Chrf score std 29.94 31.11 30.53
Ter score mean 67.79 66.19 67.13
Ter score std 30.20 31.39 30.79
Chrfpp score mean 32.76 34.36 33.42
Chrfpp score std 29.94 31.11 30.53
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 26.08 26.55 26.47
Bleu score std 25.69 26.49 25.63
Bleu 1gram precision mean 26.17 26.64 26.56
Bleu 1gram precision std 25.73 26.53 25.67
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 29.33 29.36 29.31
Bleu ref len std 14.47 14.48 14.52
Bleu pred len mean 29.25 29.28 29.23
Bleu pred len std 14.24 14.24 14.29
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 26.20 26.66 26.59
Chrf score std 25.72 26.52 25.66
Ter score mean 74.40 73.94 74.02
Ter score std 25.99 26.77 25.91
Chrfpp score mean 26.20 26.66 26.59
Chrfpp score std 25.72 26.52 25.66
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 27.31 28.63 27.55
Bleu score std 27.56 28.76 27.73
Bleu 1gram precision mean 27.47 28.75 27.71
Bleu 1gram precision std 27.61 28.80 27.78
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 30.61 30.47 30.57
Bleu ref len std 14.86 14.93 14.87
Bleu pred len mean 30.54 30.46 30.50
Bleu pred len std 14.81 14.83 14.82
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.04 0.02 0.04
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.04 0.03 0.04
Chrf score mean 27.52 28.83 27.76
Chrf score std 27.69 28.88 27.86
Ter score mean 73.47 72.14 73.22
Ter score std 27.92 29.14 28.10
Chrfpp score mean 27.52 28.83 27.76
Chrfpp score std 27.69 28.88 27.86
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 34.93 37.78 35.14
Bleu score std 29.70 31.33 29.69
Bleu 1gram precision mean 35.10 37.91 35.31
Bleu 1gram precision std 29.76 31.37 29.75
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 100.59 101.31 101.12
Bleu ref len std 48.91 49.80 49.06
Bleu pred len mean 100.42 101.37 100.88
Bleu pred len std 48.75 49.74 48.84
Bleu bp mean 0.99 1.00 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 35.55 38.40 35.76
Chrf score std 29.85 31.44 29.83
Ter score mean 75.21 71.97 74.95
Ter score std 34.89 36.82 34.81
Chrfpp score mean 35.55 38.40 35.76
Chrfpp score std 29.85 31.44 29.83
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 40.04 42.12 39.47
Bleu score std 31.35 32.61 31.61
Bleu 1gram precision mean 40.17 42.22 39.60
Bleu 1gram precision std 31.38 32.63 31.64
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 101.49 101.82 101.87
Bleu ref len std 50.11 50.22 50.16
Bleu pred len mean 101.56 101.99 101.88
Bleu pred len std 49.98 50.12 49.98
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 40.68 42.76 40.10
Chrf score std 31.39 32.64 31.65
Ter score mean 69.19 66.81 69.85
Ter score std 36.78 38.18 37.00
Chrfpp score mean 40.68 42.76 40.10
Chrfpp score std 31.39 32.64 31.65
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 40.80 43.88 41.30
Bleu score std 32.29 33.65 32.31
Bleu 1gram precision mean 40.90 43.98 41.41
Bleu 1gram precision std 32.30 33.67 32.32
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 102.43 102.86 102.68
Bleu ref len std 49.81 50.12 49.94
Bleu pred len mean 102.54 102.98 102.72
Bleu pred len std 49.77 50.05 49.84
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 41.40 44.46 41.89
Chrf score std 32.27 33.61 32.28
Ter score mean 68.32 64.76 67.77
Ter score std 37.85 39.39 37.85
Chrfpp score mean 41.40 44.46 41.89
Chrfpp score std 32.27 33.61 32.28
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 33.85 36.13 33.74
Bleu score std 28.54 30.46 28.57
Bleu 1gram precision mean 33.98 36.23 33.87
Bleu 1gram precision std 28.57 30.48 28.60
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 100.78 101.15 101.19
Bleu ref len std 49.88 49.82 49.92
Bleu pred len mean 100.70 101.22 101.07
Bleu pred len std 49.64 49.64 49.61
Bleu bp mean 0.99 1.00 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 34.48 36.74 34.37
Chrf score std 28.63 30.53 28.66
Ter score mean 76.35 73.71 76.53
Ter score std 33.54 35.72 33.57
Chrfpp score mean 34.48 36.74 34.37
Chrfpp score std 28.63 30.53 28.66
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 33.05 35.50 33.24
Bleu score std 28.46 30.64 28.72
Bleu 1gram precision mean 33.18 35.62 33.38
Bleu 1gram precision std 28.50 30.67 28.75
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 102.37 102.52 102.66
Bleu ref len std 49.92 50.07 49.96
Bleu pred len mean 102.37 102.62 102.62
Bleu pred len std 49.76 49.92 49.77
Bleu bp mean 0.99 1.00 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 33.68 36.12 33.86
Chrf score std 28.56 30.71 28.81
Ter score mean 77.43 74.65 77.20
Ter score std 33.50 36.00 33.75
Chrfpp score mean 33.68 36.12 33.86
Chrfpp score std 28.56 30.71 28.81
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 6.19 8.61 6.72
Bleu score std 12.90 16.15 13.59
Bleu 1gram precision mean 6.26 8.67 6.79
Bleu 1gram precision std 12.94 16.17 13.64
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 36.22 38.30 36.51
Bleu ref len std 20.17 20.98 20.20
Bleu pred len mean 35.45 37.76 35.74
Bleu pred len std 19.73 20.62 19.75
Bleu bp mean 0.98 0.98 0.98
Bleu bp std 0.03 0.03 0.03
Bleu ratio mean 0.98 0.99 0.98
Bleu ratio std 0.04 0.04 0.04
Chrf score mean 6.22 8.64 6.75
Chrf score std 12.93 16.17 13.62
Ter score mean 94.32 91.91 93.79
Ter score std 13.12 16.39 13.81
Chrfpp score mean 6.22 8.64 6.75
Chrfpp score std 12.93 16.17 13.62
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 14.56 15.30 14.61
Bleu score std 20.15 20.98 19.70
Bleu 1gram precision mean 14.71 15.43 14.76
Bleu 1gram precision std 20.30 21.10 19.86
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 40.61 40.93 40.74
Bleu ref len std 19.85 20.05 19.91
Bleu pred len mean 40.24 40.66 40.37
Bleu pred len std 19.70 19.97 19.76
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 0.99 0.99 0.99
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 14.65 15.39 14.70
Chrf score std 20.22 21.05 19.77
Ter score mean 86.02 85.28 85.97
Ter score std 20.37 21.21 19.92
Chrfpp score mean 14.65 15.39 14.70
Chrfpp score std 20.22 21.05 19.77

### v3

@@ Appendix C Additional Experiment Results
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 17.29 19.21 18.31 19.69
Bleu score std 22.88 24.37 24.26 25.91
Bleu 1gram precision mean 17.37 19.30 18.40 19.74
Bleu 1gram precision std 22.94 24.43 24.32 25.96
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 24.74 26.52 24.75 26.10
Bleu ref len std 12.69 13.49 12.72 13.40
Bleu pred len mean 24.59 26.35 24.60 26.09
Bleu pred len std 12.46 13.22 12.49 13.29
Bleu bp mean 0.99 0.99 0.99 1.00
Bleu bp std 0.02 0.02 0.02 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03 0.03
Chrf score mean 17.45 19.36 18.47 19.84
Chrf score std 23.08 24.54 24.43 26.07
Ter score mean 83.26 81.33 82.23 80.86
Ter score std 23.01 24.53 24.40 26.07
Chrfpp score mean 17.45 19.36 18.47 19.84
Chrfpp score std 23.08 24.54 24.43 26.07
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 26.26 27.23 27.06 29.01
Bleu score std 24.84 26.18 25.59 27.21
Bleu 1gram precision mean 26.51 27.50 27.32 29.21
Bleu 1gram precision std 24.89 26.23 25.65 27.26
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 28.70 29.35 28.76 29.13
Bleu ref len std 14.62 14.70 14.65 14.61
Bleu pred len mean 28.24 28.86 28.30 28.85
Bleu pred len std 13.91 13.97 13.96 14.16
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.03 0.03 0.03 0.03
Bleu ratio mean 0.99 0.99 0.99 1.00
Bleu ratio std 0.04 0.04 0.04 0.04
Chrf score mean 26.43 27.41 27.23 29.17
Chrf score std 24.92 26.25 25.66 27.27
Ter score mean 74.32 73.34 73.51 71.57
Ter score std 25.21 26.56 25.97 27.60
Chrfpp score mean 26.43 27.41 27.23 29.17
Chrfpp score std 24.92 26.25 25.66 27.27
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 32.64 34.24 33.30 34.12
Bleu score std 29.93 31.10 30.52 30.34
Bleu 1gram precision mean 32.69 34.30 33.34 34.16
Bleu 1gram precision std 29.94 31.12 30.53 30.35
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 29.97 30.02 29.94 29.97
Bleu ref len std 14.74 14.79 14.77 14.75
Bleu pred len mean 30.02 30.06 30.00 30.02
Bleu pred len std 14.73 14.74 14.77 14.72
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 32.76 34.36 33.42 34.24
Chrf score std 29.94 31.11 30.53 30.35
Ter score mean 67.79 66.19 67.13 66.31
Ter score std 30.20 31.39 30.79 30.62
Chrfpp score mean 32.76 34.36 33.42 34.24
Chrfpp score std 29.94 31.11 30.53 30.35
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 26.08 26.55 26.47 26.53
Bleu score std 25.69 26.49 25.63 26.37
Bleu 1gram precision mean 26.17 26.64 26.56 26.62
Bleu 1gram precision std 25.73 26.53 25.67 26.40
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 29.33 29.36 29.31 29.37
Bleu ref len std 14.47 14.48 14.52 14.49
Bleu pred len mean 29.25 29.28 29.23 29.28
Bleu pred len std 14.24 14.24 14.29 14.22
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.02 0.02 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03 0.03
Chrf score mean 26.20 26.66 26.59 26.65
Chrf score std 25.72 26.52 25.66 26.40
Ter score mean 74.40 73.94 74.02 73.95
Ter score std 25.99 26.77 25.91 26.64
Chrfpp score mean 26.20 26.66 26.59 26.65
Chrfpp score std 25.72 26.52 25.66 26.40
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 27.31 28.63 27.55 27.84
Bleu score std 27.56 28.76 27.73 28.04
Bleu 1gram precision mean 27.47 28.75 27.71 27.96
Bleu 1gram precision std 27.61 28.80 27.78 28.10
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 30.61 30.47 30.57 30.51
Bleu ref len std 14.86 14.93 14.87 14.98
Bleu pred len mean 30.54 30.46 30.50 30.51
Bleu pred len std 14.81 14.83 14.82 14.88
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.04 0.02 0.04 0.02
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.04 0.03 0.04 0.03
Chrf score mean 27.52 28.83 27.76 28.04
Chrf score std 27.69 28.88 27.86 28.16
Ter score mean 73.47 72.14 73.22 72.94
Ter score std 27.92 29.14 28.10 28.42
Chrfpp score mean 27.52 28.83 27.76 28.04
Chrfpp score std 27.69 28.88 27.86 28.16
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 34.93 37.78 35.14 38.16
Bleu score std 29.70 31.33 29.69 31.98
Bleu 1gram precision mean 35.10 37.91 35.31 38.28
Bleu 1gram precision std 29.76 31.37 29.75 32.02
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 100.59 101.31 101.12 101.26
Bleu ref len std 48.91 49.80 49.06 49.16
Bleu pred len mean 100.42 101.37 100.88 101.26
Bleu pred len std 48.75 49.74 48.84 49.02
Bleu bp mean 0.99 1.00 0.99 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 35.55 38.40 35.76 38.76
Chrf score std 29.85 31.44 29.83 32.08
Ter score mean 75.21 71.97 74.95 71.49
Ter score std 34.89 36.82 34.81 37.49
Chrfpp score mean 35.55 38.40 35.76 38.76
Chrfpp score std 29.85 31.44 29.83 32.08
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 40.04 42.12 39.47 41.72
Bleu score std 31.35 32.61 31.61 32.25
Bleu 1gram precision mean 40.17 42.22 39.60 41.84
Bleu 1gram precision std 31.38 32.63 31.64 32.29
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 101.49 101.82 101.87 102.01
Bleu ref len std 50.11 50.22 50.16 50.25
Bleu pred len mean 101.56 101.99 101.88 102.13
Bleu pred len std 49.98 50.12 49.98 50.11
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 40.68 42.76 40.10 42.35
Chrf score std 31.39 32.64 31.65 32.26
Ter score mean 69.19 66.81 69.85 67.29
Ter score std 36.78 38.18 37.00 37.78
Chrfpp score mean 40.68 42.76 40.10 42.35
Chrfpp score std 31.39 32.64 31.65 32.26
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 40.80 43.88 41.30 43.18
Bleu score std 32.29 33.65 32.31 33.31
Bleu 1gram precision mean 40.90 43.98 41.41 43.27
Bleu 1gram precision std 32.30 33.67 32.32 33.33
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 102.43 102.86 102.68 102.85
Bleu ref len std 49.81 50.12 49.94 50.07
Bleu pred len mean 102.54 102.98 102.72 102.97
Bleu pred len std 49.77 50.05 49.84 49.98
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 41.40 44.46 41.89 43.77
Chrf score std 32.27 33.61 32.28 33.28
Ter score mean 68.32 64.76 67.77 65.56
Ter score std 37.85 39.39 37.85 39.02
Chrfpp score mean 41.40 44.46 41.89 43.77
Chrfpp score std 32.27 33.61 32.28 33.28
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 33.85 36.13 33.74 36.33
Bleu score std 28.54 30.46 28.57 30.77
Bleu 1gram precision mean 33.98 36.23 33.87 36.44
Bleu 1gram precision std 28.57 30.48 28.60 30.79
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 100.78 101.15 101.19 101.29
Bleu ref len std 49.88 49.82 49.92 49.92
Bleu pred len mean 100.70 101.22 101.07 101.32
Bleu pred len std 49.64 49.64 49.61 49.73
Bleu bp mean 0.99 1.00 0.99 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 34.48 36.74 34.37 36.94
Chrf score std 28.63 30.53 28.66 30.82
Ter score mean 76.35 73.71 76.53 73.46
Ter score std 33.54 35.72 33.57 36.11
Chrfpp score mean 34.48 36.74 34.37 36.94
Chrfpp score std 28.63 30.53 28.66 30.82
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 33.05 35.50 33.24 35.02
Bleu score std 28.46 30.64 28.72 30.28
Bleu 1gram precision mean 33.18 35.62 33.38 35.15
Bleu 1gram precision std 28.50 30.67 28.75 30.32
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 102.37 102.52 102.66 102.71
Bleu ref len std 49.92 50.07 49.96 50.03
Bleu pred len mean 102.37 102.62 102.62 102.77
Bleu pred len std 49.76 49.92 49.77 49.85
Bleu bp mean 0.99 1.00 0.99 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 33.68 36.12 33.86 35.64
Chrf score std 28.56 30.71 28.81 30.36
Ter score mean 77.43 74.65 77.20 75.15
Ter score std 33.50 36.00 33.75 35.54
Chrfpp score mean 33.68 36.12 33.86 35.64
Chrfpp score std 28.56 30.71 28.81 30.36
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 6.19 8.61 6.72 9.14
Bleu score std 12.90 16.15 13.59 17.19
Bleu 1gram precision mean 6.26 8.67 6.79 9.18
Bleu 1gram precision std 12.94 16.17 13.64 17.20
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 36.22 38.30 36.51 38.18
Bleu ref len std 20.17 20.98 20.20 20.76
Bleu pred len mean 35.45 37.76 35.74 37.85
Bleu pred len std 19.73 20.62 19.75 20.74
Bleu bp mean 0.98 0.98 0.98 0.99
Bleu bp std 0.03 0.03 0.03 0.03
Bleu ratio mean 0.98 0.99 0.98 0.99
Bleu ratio std 0.04 0.04 0.04 0.03
Chrf score mean 6.22 8.64 6.75 9.17
Chrf score std 12.93 16.17 13.62 17.20
Ter score mean 94.32 91.91 93.79 91.38
Ter score std 13.12 16.39 13.81 17.42
Chrfpp score mean 6.22 8.64 6.75 9.17
Chrfpp score std 12.93 16.17 13.62 17.20
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 14.56 15.30 14.61 16.94
Bleu score std 20.15 20.98 19.70 22.63
Bleu 1gram precision mean 14.71 15.43 14.76 17.02
Bleu 1gram precision std 20.30 21.10 19.86 22.68
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 40.61 40.93 40.74 41.08
Bleu ref len std 19.85 20.05 19.91 20.14
Bleu pred len mean 40.24 40.66 40.37 40.98
Bleu pred len std 19.70 19.97 19.76 20.10
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.02 0.02 0.02 0.02
Bleu ratio mean 0.99 0.99 0.99 1.00
Bleu ratio std 0.03 0.03 0.03 0.02
Chrf score mean 14.65 15.39 14.70 17.02
Chrf score std 20.22 21.05 19.77 22.67
Ter score mean 86.02 85.28 85.97 83.65
Ter score std 20.37 21.21 19.92 22.89
Chrfpp score mean 14.65 15.39 14.70 17.02
Chrfpp score std 20.22 21.05 19.77 22.67